OpenStudy (mhchen):

Multiplying variances: Suppose a variance = 2. If you were to multiply it by 4, how would it transform it. Is it just directly multiplying it by 4, or do you have to square the 4 first? I know you can't add variances by a constant, idk if you directly multiply them. Formal Notation Shown Below:

OpenStudy (mhchen):

\[\sigma ^{2}_{x} = 2~~~~~~~\sigma^{2}_{4x}=?\]

OpenStudy (amorfide):

If Var(X) = 2.5 then, Var(2X) = 2^2 × Var(X) = 4 × 2.5 = 10 Var(4X − 3) = 4^2 × Var(X) = 16 × 2.5 = 40 so to my understanding you would do 4^2 x 2

OpenStudy (mhchen):

That's what I thought as well. ty.